OpenStudy (anonymous):

write an equation in slope intercept form with slope 8 and containing the point (5,3)

OpenStudy (anonymous):

You use y-y1=m(x-x1) for this just plug in the points and the slope c:

OpenStudy (anonymous):

so it would be y-3=8(x-5) distribute and you get y=8x-37

OpenStudy (anonymous):

thanks ! @Adri1580 what about ... graph the equation y=-2/3x+4

OpenStudy (anonymous):

First you move up four on the y axis. So your on the point (0,4). Draw a dot there c: After you move down 2 points and 3 to the right. so you would be at the point (3,2). Draw a dot on this point as well and now you just draw a line through these two points and you have your graph!

OpenStudy (anonymous):

ok thanks .. also what about write an equation for the line containing the points (-2,3) and (3,-4) @Adri1580

OpenStudy (anonymous):

 my sucky drawing xD

OpenStudy (anonymous):

For this you must use the slope formula to first find the slope. (y2-y1)/(x2-x1). The slope for this equation would be -7/5. Now we just do what we did with the other equation with this one! y=-7/5x+1/5

OpenStudy (anonymous):

thank you !!! ok whatabout find the slope of the line containing (-7,-3) (-5,3) @Adri1580

OpenStudy (anonymous):

and also if you dont mind find the slope and y intercept. 3x-8y=-31 @Adri1580 do i just divide them by 8 ?

OpenStudy (anonymous):

SORRY okay the slope for that line would be -6/12 if you use the slope formula (y2-y1)/(x2-x1). You would have to change the equation into slope intercept form so you would have y=3/8x+31/8 the slope would be 3/8 and the y intercept would be (0, 31/8)

OpenStudy (anonymous):

sorry it took so long my computer froze >.<

OpenStudy (anonymous):

Oh wait the slope for the first question isnt -6/12 its 6/2